OpenStudy (anonymous):

What are the solutions of 6x2 − x − 15 = 0? x = 5 over 6, x = −3 x = −5 over 6, x = 3 x = 3 over 2, x = −5 over 3 x = −3 over 2, x =5 over 3

OpenStudy (shamil98):

Use the quadratic formula. \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] ax^2 + bx + c = 0 so your a b and c values are?